Car travels 72km/hr in 5 sec find acceleration and distance travelled? If car takes 2.5 seconds to stop calculate the new acceleration and find the distance a car takes to stop
1
Expert's answer
2020-08-12T16:18:58-0400

"v_0=0"; "72(km\/hr)=20(m\/s)"


(a)


"v=at\to a=v/t=20/5=4(m/s^2)"


"S=at^2/2=4\cdot5^2/2=50(m)"


(b)


"v_0=20(m\/s)"


"v=v_0-at\to a=v_0/t=20/2.5=8(m/s^2)"


"S=\frac{v^2-v_0^2}{2a}=\frac{20^2}{2\cdot8}=25(m)"